# EW corrections and Heavy Boson Radiation at a high-energy muon collider

### v1

@@ 4.1.2 Importance of logarithms involving ratios of invariants
Figure 7: Same as Fig. 6, for the $H$ boson $p_{T}$ distribution in $\mu^{+}\mu^{-}\to ZH$.
The case of $p_{T}(H)$ distribution in $ZH$ production, Fig. 7, shows a very similar pattern at the differential level, although the impact of $\delta_{s\rightarrow r_{kl}}$ is smaller. However, since the $p_{T}(H)$ distribution is much less flat, at the inclusive level also with $\delta_{s\rightarrow r_{kl}}=0$ a good accuracy for the ${\rm SDK_{weak}}$ prediction is present. It is interesting to note that for small values of $p_{T}(H)$ at 10 TeV the case with $\delta_{s\rightarrow r_{kl}}=0$ yields smaller values of $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$. As said at the beginning of the previous section, this is not per se a sign of better accuracy. Indeed this effect is due to missing logarithms among invariants that accidentally compensate the large non-logarithmically enhanced ${\cal O}(\alpha)$ component already discussed in the case of Fig. 5.
@@
To the best of our understanding, this $10\%$ discrepancy is not due to a large non-logarithmically enhanced ${\cal O}(\alpha)$ component, as in the case of $ZH$ distributions, but to large logarithms that the ${\rm SDK_{weak}}$ scheme is not able to capture even retaining the $\delta_{s\rightarrow r_{kl}}$ term. This can better understood by looking at the $p_{T}(Z_{3})$ distribution, the $p_{T}$ of the softest $Z$ boson, in Fig. 9 and the $\Delta R(Z_{1},Z_{2})$ distribution in Fig. 10, in particular at 10 TeV. First we observe that for very large $p_{T}(Z_{3})$, meaning all $Z$ bosons that are hard and so all the invariants that are large, $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\to 0$, while the same quantity constantly grows in the opposite direction. Second, for $\Delta R(Z_{1},Z_{2})\lesssim\pi$, $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 0$, while for $\Delta R(Z_{1},Z_{2})\gtrsim\pi$ the same quantity jumps to $\sim 10\%$ and remains constant up to $\Delta R(Z_{1},Z_{2})\simeq 5$. The dominant kinematic configuration is $Z_{1}$ and $Z_{2}$ that are almost back-to-back, i.e. $\Delta R(Z_{1},Z_{2})\simeq\pi$. The region $\Delta R(Z_{1},Z_{2})\lesssim\pi$ is dominated by large values of $p_{T}(Z_{3})$, which therefore are correlated and both show $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 0$. Instead, in the region $\Delta R(Z_{1},Z_{2})\gtrsim\pi$ large contributions from the $\rm NLO_{EW}$ prediction originate from the $ZZZ\gamma$ final state, which allows a further recoil and enhances $|\eta(Z_{1})-\eta(Z_{2})|$ and in turn $\Delta R(Z_{1},Z_{2})$. This dynamics is only captured by the $\rm NLO_{EW}$ prediction. Another peculiar behaviour is observed in the first bins of the distribution, where the EWSL prediction departs from the NLO EW. This is due to the fact that, in a Born-like kinematics, the first bin can be filled only when all $Z$ bosons have equal transverse momentum. Photon radiation, captured only by the NLO EW prediction, lifts such a constraint, and can thus enhance this region All in all, for this process, the inclusion of $\delta_{s\rightarrow r_{kl}}$ is crucial for improving the approximation. Nevertheless, non-negligible effects are not captured.
@@
As clearly stated in Refs. Denner:2000jv ; Denner:2001gw , the DP algorithm assumes that the helicity configuration considered is not mass suppressed, i.e., as said in Sec. 2.2.1, that it scales as ${\cal{M}}\propto s^{\frac{2-n}{2}}$ for a $2\to n$ process. In general, in the SM, at least one of the helicity configurations of the processes considered is typically not mass suppressed and therefore at high energies is very enhanced w.r.t. the other ones. [^17] This enhancement is very large, by at least ${\cal O}(s/M_{W}^{2})$. For this reason, even blindly applying the DP algorithm to all helicity configurations, regardless if they are or are not mass-suppressed, the prediction for the sum over the polarisations is consistent, namely it corresponds to the high-energy limit $M_{W}^{2}/s\to 0$. In other words, even if the algorithm returns wrong results for the mass-suppressed helicity configurations, they are so suppressed that the relative impact in the sum over the helicity configurations is completely negligible. However, it is known that there can be processes where none of the helicity configurations are not mass-suppressed, such as Higgs VBF production. [^18] In an upcoming paper, this aspect will be discussed in detail in the context of the SM Effective-Filed theory (SMEFT) elfaham:xxx , where these cases are much more common. For such cases, the DP algorithm is known to no be working.
@@
Figure 13: Same as Fig. 12, for the $Z$ boson $p_{T}$ distribution in $\mu^{+}\mu^{-}\to ZHH$.
@@ 4.2 Beyond NLO EW: the relevance of resummation
Figure 15: Same as Fig. 14, for $p_{T}(Z_{2})$ distribution in $\mu^{+}\mu^{-}\to ZZ$
@@
Figure 16: Same as Fig. 14, for the Higgs boson $p_{T}$ distribution in $\mu^{+}\mu^{-}\to ZH$
@@ 4.3.1 Final states: $t\bar{t}$, $W^{+}W^{-}$
A couple of features are manifest in Fig. 18. First, the contribution of $t\bar{t}Z$ is much larger than the one from $t\bar{t}H$, so $\delta_{\rm HBR}$ is dominated by the former. Second, we see that at very large $p_{T}(t)$, if top-quarks are recombined with the $B=H,V$ the contribution from both $t\bar{t}H$ and $t\bar{t}Z$ is much larger (solid vs. dashed). Indeed, only via the recombination of a top with the HBR the $t\bar{t}$ pair can reach the value $m(t\bar{t})\simeq s\simeq S$, which would not be possible otherwise even with soft HBR since $M_{B}\neq 0$. We will come back with more details on this point later in the discussion. Third, even considering such a recombination procedure, the total HBR contribution is negligible w.r.t. the effects of the same perturbative order induced by NLO EW corrections, both at 3 and 10 TeV. This can be seen by comparing the absolute values of $\delta_{\rm HBR}$ and $\delta_{\rm NLO_{EW}}$ over the spectrum for $p_{T}(t)\gtrsim 250~{}\,\textrm{GeV}$ ($p_{T}(t)\gtrsim 800~{}\,\textrm{GeV}$) at 3 TeV (10 TeV) collisions. Only for smaller values of $p_{T}(t)$ the HBR contribution is very large, similar to the case of the $\rm NLO_{EW}$ predictions. The same argument presented for explaining the large contribution from the real photon radiation and in turn the $\rm NLO_{EW}$ prediction in the discussion of Fig. 1 can be repeated here for the HBR. Finally, we see that, as a consequence of the previous point, $\delta_{\rm NLO_{EW}}\simeq\delta_{\rm NLO_{EW}+HBR}$ for $p_{T}(t)\gtrsim 250~{}\,\textrm{GeV}$ ($p_{T}(t)\gtrsim 800~{}\,\textrm{GeV}$) at 3 TeV (10 TeV) collisions. This implies the EWSL from weak virtual corrections, i.e. the $\delta_{{\rm SDK_{weak}}}$ prediction which we have been shown to be the dominant component of $\delta_{\rm NLO_{EW}}$ in the discussion of Fig 1, are minimally compensated by those from real radiation.
@@
Figure 21: Same as Fig. 19, but for the $p_{T}(W^{+})$ distribution in $\mu^{+}\mu^{-}\to W^{+}W^{-}$. Also in this case $m(t\bar{t})>0.5\sqrt{S}$.
@@ 4.3.2 EW jets in multi-$V$ production
We now consider the case of inclusive EW-dijet production, $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, where in the EW-jet $j_{\rm EW}$ we cluster $W$ and $Z$ bosons. In other words, the LO originates from the processes $\mu^{+}\mu^{-}\to VV$ with $V=W,Z$, which correspond to the final states $WW$ and $ZZ$. Having found in the previous section that the impact of the Higgs boson radiation is minimal, we consider as a contribution to the HBR only the processes involving an additional $V=W,Z$ radiation, $\mu^{+}\mu^{-}\to 3V$, which correspond to the final states $ZZZ$ and $WWZ$. Clearly, the (at least) two $j_{\rm EW}$’s that are required may be also associated with $W$ and $Z$, respectively, or also with a $WZ$ pair and a $Z$ respectively. Several combinations are possible and some of them, as the ones just mentioned and unlike the cases discussed in the previous section, do not feature the same “partonic” final state of the LO, which can feature only $ZZ$ and $WW$. We will consider also the case of a double HBR and therefore the $4V$ final states: $ZZZZ$, $WWWW$ and $WWZZ$ [^23] As a side comment, such processes have been shown to be very sensitive to possible anomalous interactions of the Higgs boson with the muon Celada:2023oji ; Han:2021lnp and therefore are of particular relevance in the muon collider physics program.
Figure 23: The $p_{T}(j_{{\rm EW},1})$ distribution in $\mu^{+}\mu^{-}\to 2j_{\rm EW}$. The left (right) plot shows results at $\sqrt{S}=3$ TeV ($\sqrt{S}=10$ TeV). The histograms show LO (dashed) and NLO (solid predictions) for $W^{+}W^{-}$ (green), $ZZ$ (blue), $2V=W^{+}W^{-}+ZZ$ (orange), $3V$ (red), and $4V$ (violet, only at the LO). In the second inset, the quantities $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$ and $\delta_{\rm NLO_{EW}+HBR}$ are shown respectively as black-solid and black-dashed lines. These quantities are defined by Eq. (3.26) in terms of respectively Eq. (3.25) and Eq. (3.24). The solid grey line shows the quantity $\delta_{{\rm SDK_{weak}}}^{2}/2$, which enters only in $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$.
@@ 5 Conclusions and outlook
We have performed a comprehensive study of EW corrections at a future multi-TeV muon collider, using the MadGraph5_aMC@NLO framework. The focus of our work has been on direct-production processes. For these colliders, EW corrections are sizeable, and their inclusion is mandatory not just for precision studies, but even for proper estimates of scattering rates. Within MG5_aMC, EW corrections can be computed either exactly, or in the Sudakov (high-energy) approximation, exploiting the DP algorithm. The latter case has clear advantages in terms of speed and numerical stability, however its accuracy must always be validated against the exact NLO EW computation. Indeed, in this respect, we have shown the benefit of employing a purely-weak version of the DP approach, and the importance of including extra angular-dependent terms. Both aspects were first highlighted in Ref. Pagani:2021vyk . Still, we have discussed exceptional cases when the high-energy approximation as obtained from the DP method fails, sometimes in a rather spectacular way, as in the case of $ZHH$ production.
Besides, we have discussed when, due to their large size, the resummation of EW corrections is necessary to have reliable estimates of rates. Although based on an approximate formula, our findings show that resummation is mandatory for very high collider energies or for multi-boson processes or, in general whenever EW corrections easily exceed -50% with respect to the LO rate. In other cases, resummation is still beneficial for precision studies.
@@
Acknowledgments
Computational resources have been provided by the supercomputing facilities of the Université catholique de Louvain (CISM/UCL) and the Consortium des Équipements de Calcul Intensif en Fédération Wallonie Bruxelles (CÉCI) funded by the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS) under convention 2.5020.11 and by the Walloon Region. DP acknowledges support from the DESY computing infrastructures provided through the DESY theory group for the initial stages of this work. The work of M.Z. has been partly supported by the “Programma per Giovani Ricercatori Rita Levi Montalcini” granted by the Italian Ministero dell’Università e della Ricerca (MUR). We also acknowledge support from the COMETA COST Action CA22130.

### v2

@@ 4.1.2 Importance of logarithms involving ratios of invariants
Figure 7: Same as Fig. 6, for the $p_{T}(H)$ distribution in $\mu^{+}\mu^{-}\to ZH$.
The case of $p_{T}(H)$ distribution in $ZH$ production, Fig. 7, shows a very similar pattern at the differential level, although the impact of $\delta_{s\rightarrow r_{kl}}$ is smaller. Moreover, since the $p_{T}(H)$ distribution is much less flat, it is manifest that at the inclusive level the accuracy for the ${\rm SDK_{weak}}$ prediction is not affected by the assumption $\delta_{s\rightarrow r_{kl}}=0$. It is interesting to note that for small values of $p_{T}(H)$ at 10 TeV the case with $\delta_{s\rightarrow r_{kl}}=0$ yields smaller values of $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}$. As said at the beginning of the previous section, this is not per se a sign of better accuracy. Indeed this effect is due to missing logarithms among invariants that accidentally compensate the large non-logarithmically enhanced ${\cal O}(\alpha)$ component already discussed in the case of Fig. 5.
@@
To the best of our understanding, this $10\%$ discrepancy is not due to a large non-logarithmically enhanced ${\cal O}(\alpha)$ component, as in the case of $ZH$ distributions, but to large logarithms that the ${\rm SDK_{weak}}$ scheme is not able to capture even retaining the $\delta_{s\rightarrow r_{kl}}$ term. This can be better understood by looking at the $p_{T}(Z_{3})$ distribution, the $p_{T}$ of the softest $Z$ boson, in Fig. 9 and the $\Delta R(Z_{1},Z_{2})$ distribution in Fig. 10, in particular at 10 TeV. First we observe that for very large $p_{T}(Z_{3})$, meaning all $Z$ bosons that are hard and so all the invariants that are large, $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\to 0$, while the same quantity constantly grows in the opposite direction. Second, for $\Delta R(Z_{1},Z_{2})\lesssim\pi$, $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 0$, while for $\Delta R(Z_{1},Z_{2})\gtrsim\pi$ the same quantity jumps to $\sim 10\%$ and remains constant up to $\Delta R(Z_{1},Z_{2})\simeq 5$. The dominant kinematic configuration is $Z_{1}$ and $Z_{2}$ that are almost back-to-back, i.e. $\Delta R(Z_{1},Z_{2})\simeq\pi$. The region $\Delta R(Z_{1},Z_{2})\lesssim\pi$ is dominated by large values of $p_{T}(Z_{3})$, which therefore are correlated and both show $\delta_{{\rm SDK_{weak}}}-\delta_{\rm NLO_{EW}}\simeq 0$. Instead, in the region $\Delta R(Z_{1},Z_{2})\gtrsim\pi$ large contributions from the $\rm NLO_{EW}$ prediction originate from the $ZZZ\gamma$ final state, which allows a further recoil and enhances $|\eta(Z_{1})-\eta(Z_{2})|$ and in turn $\Delta R(Z_{1},Z_{2})$. This dynamics is only captured by the $\rm NLO_{EW}$ prediction. Another peculiar behaviour is observed in the first bins of the distribution, where the EWSL prediction departs from the NLO EW. This is due to the fact that, in a Born-like kinematics, the first bin can be filled only when all $Z$ bosons have equal transverse momentum. Photon radiation, captured only by the NLO EW prediction, lifts such a constraint, and can thus enhance this region All in all, for this process, the inclusion of $\delta_{s\rightarrow r_{kl}}$ is crucial for improving the approximation. Nevertheless, non-negligible effects are not captured.
@@
As clearly stated in Refs. Denner:2000jv ; Denner:2001gw , the DP algorithm assumes that the helicity configuration considered is not mass suppressed, i.e., as said in Sec. 2.2.1, that it scales as ${\cal{M}}\propto s^{\frac{2-n}{2}}$ for a $2\to n$ process. In general, in the SM, at least one of the helicity configurations of the processes considered is typically not mass suppressed and therefore at high energies is very enhanced w.r.t. the other ones. [^18] This enhancement is very large, by at least ${\cal O}(s/M_{W}^{2})$. For this reason, even blindly applying the DP algorithm to all helicity configurations, regardless if they are or are not mass-suppressed, the prediction for the sum over the polarisations is consistent, namely it corresponds to the high-energy limit $M_{W}^{2}/s\to 0$. In other words, even if the algorithm returns wrong results for the mass-suppressed helicity configurations, they are so suppressed that the relative impact in the sum over the helicity configurations is completely negligible. However, it is known that there can be processes where none of the helicity configurations are not mass-suppressed, such as Higgs VBF production. [^19] In an upcoming paper, this aspect will be discussed in detail in the context of the SM Effective-Filed theory (SMEFT) ElFaham:2024egs , where these cases are much more common. For such cases, the DP algorithm is known to no be working.
@@
Figure 13: Same as Fig. 12, for the $p_{T}(Z)$ distribution in $\mu^{+}\mu^{-}\to ZHH$.
@@ 4.2 Beyond NLO EW: the relevance of resummation
Figure 15: Same as Fig. 14, for the $p_{T}(Z_{2})$ distribution in $\mu^{+}\mu^{-}\to ZZ$
@@
Figure 16: Same as Fig. 14, for the $p_{T}(H)$ distribution in $\mu^{+}\mu^{-}\to ZH$
@@ 4.3.1 Final states: $t\bar{t}$, $W^{+}W^{-}$
A couple of features are manifest in Fig. 18. First, the contribution of $t\bar{t}Z$ is much larger than the one from $t\bar{t}H$, so $\delta_{\rm HBR}$ is dominated by the former. The origin of this difference is due to the fact that at high energy the soft emission of a Higgs boson is logarithmically enhanced, as in the case of the $Z$ emission, but it is also mass suppressed ($m_{t}^{2}/s$), at variance with the $Z$ emission. [^22] In the case of loop corrections, this is precisely the same reason which allows to neglect the contribution of Higgs loops in the calculation of EWSL of soft origin. Second, we see that at very large $p_{T}(t)$, if top-quarks are recombined with the $B=H,V$ the contribution from both $t\bar{t}H$ and $t\bar{t}Z$ is much larger (solid vs. dashed). Indeed, only via the recombination of a top with the HBR the $t\bar{t}$ pair can reach the value $m(t\bar{t})\simeq s\simeq S$, which would not be possible otherwise even with soft HBR since $M_{B}\neq 0$. We will come back with more details on this point later in the discussion. Third, even considering such a recombination procedure, the total HBR contribution is negligible w.r.t. the effects of the same perturbative order induced by NLO EW corrections, both at 3 and 10 TeV. This can be seen by comparing the absolute values of $\delta_{\rm HBR}$ and $\delta_{\rm NLO_{EW}}$ over the spectrum for $p_{T}(t)\gtrsim 250~{}\,\textrm{GeV}$ ($p_{T}(t)\gtrsim 800~{}\,\textrm{GeV}$) at 3 TeV (10 TeV) collisions. Only for smaller values of $p_{T}(t)$ the HBR contribution is very large, similar to the case of the $\rm NLO_{EW}$ predictions. The same argument presented for explaining the large contribution from the real photon radiation and in turn the $\rm NLO_{EW}$ prediction in the discussion of Fig. 1 can be repeated here for the HBR. Finally, we see that, as a consequence of the previous point, $\delta_{\rm NLO_{EW}}\simeq\delta_{\rm NLO_{EW}+HBR}$ for $p_{T}(t)\gtrsim 250~{}\,\textrm{GeV}$ ($p_{T}(t)\gtrsim 800~{}\,\textrm{GeV}$) at 3 TeV (10 TeV) collisions. This implies the EWSL from weak virtual corrections, i.e. the $\delta_{{\rm SDK_{weak}}}$ prediction which we have been shown to be the dominant component of $\delta_{\rm NLO_{EW}}$ in the discussion of Fig 1, are minimally compensated by those from real radiation.
@@
Figure 21: Same as Fig. 19, but for the $p_{T}(W^{+})$ distribution in $\mu^{+}\mu^{-}\to W^{+}W^{-}$. Also in this case $m(W^{+}W^{-})>0.5\sqrt{S}$.
@@ 4.3.2 EW jets in multi-$V$ production
We now consider the case of inclusive EW-dijet production, $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, where in the EW-jet $j_{\rm EW}$ we cluster $W$ and $Z$ bosons. In other words, the LO originates from the processes $\mu^{+}\mu^{-}\to VV$ with $V=W,Z$, which correspond to the final states $WW$ and $ZZ$. Having found in the previous section that the impact of the Higgs boson radiation is minimal, we consider as a contribution to the HBR only the processes involving an additional $V=W,Z$ radiation, $\mu^{+}\mu^{-}\to 3V$, which correspond to the final states $ZZZ$ and $WWZ$. Clearly, the (at least) two $j_{\rm EW}$’s that are required may be also associated with $W$ and $Z$, respectively, or also with a $WZ$ pair and a $W$ respectively. Several combinations are possible and some of them, as the ones just mentioned and unlike the cases discussed in the previous section, do not feature the same “partonic” final state of the LO, which can feature only $ZZ$ and $WW$. We will consider also the case of a double HBR and therefore the $4V$ final states: $ZZZZ$, $WWWW$ and $WWZZ$ [^25] As a side comment, such processes have been shown to be very sensitive to possible anomalous interactions of the Higgs boson with the muon Celada:2023oji ; Han:2021lnp and therefore are of particular relevance in the muon collider physics program.
Figure 23: The $p_{T}(j_{{\rm EW},1})$ distribution in $\mu^{+}\mu^{-}\to 2j_{\rm EW}$, with $m(2j_{\rm EW})>0.8\sqrt{S}$. The left (right) plot shows results at $\sqrt{S}=3$ TeV ($\sqrt{S}=10$ TeV). The histograms show LO (dashed) and NLO (solid predictions) for $W^{+}W^{-}$ (green), $ZZ$ (blue), $2V=W^{+}W^{-}+ZZ$ (orange), $3V$ (red), and $4V$ (violet, only at the LO). In the second inset, the quantities $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$ and $\delta_{\rm NLO_{EW}+HBR}$ are shown respectively as black-solid and black-dashed lines. These quantities are defined by Eq. (3.26) in terms of respectively Eq. (3.25) and Eq. (3.24). The solid grey line shows the quantity $\delta_{{\rm SDK_{weak}}}^{2}/2$, which enters only in $\delta_{\rm nNLO_{EW}+HBR_{\rm NLO}}$.
@@ 5 Conclusions and outlook
We have performed a comprehensive study of NLO EW corrections at a future multi-TeV muon collider, using the MadGraph5_aMC@NLO framework. The focus of our work has been on direct-production processes. For these colliders, NLO EW corrections are sizeable, and their inclusion is mandatory not just for precision studies, but even for proper estimates of scattering rates. Within MG5_aMC, EW corrections can be computed either exactly at NLO, or in the Sudakov (high-energy) approximation, exploiting the DP algorithm. The latter case has clear advantages in terms of speed and numerical stability, however its accuracy must always be validated against the exact NLO EW computation. Indeed, in this respect, we have shown the benefit of employing a purely-weak version of the DP approach, and the importance of including extra angular-dependent terms. Both aspects were first highlighted in Ref. Pagani:2021vyk . Still, we have discussed exceptional cases when the high-energy approximation as obtained from the DP method fails, sometimes in a rather spectacular way, as in the case of $ZHH$ production.
Besides, we have discussed when, due to their large size, the resummation of EW Sudakov logarithms is necessary in order to have sensible predictions or even simply positive cross sections. Although based on an approximate formula, our findings show that resummation is mandatory for multi-boson processes at 10 TeV or in general whenever EW corrections approach -100% with respect to the LO rate. However, there are also processes, as, e.g., the top-quark production, where resummation is definitely necessary for precision studies, but its impact on top of NLO EW predictions is below the 5%(10%) level at 3(10) TeV. Thus, dedicated studies on the resummation of Sudakov logarithms for specific processes and at different energies would be desirable.
@@
Acknowledgements.
Computational resources have been provided by the supercomputing facilities of the Université catholique de Louvain (CISM/UCL) and the Consortium des Équipements de Calcul Intensif en Fédération Wallonie Bruxelles (CÉCI) funded by the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS) under convention 2.5020.11 and by the Walloon Region. YM acknowledges the support as a Postdoctoral Fellow of the Fond de la Recherche Scientifique de Belgique (F.R.S.-FNRS), Belgium. DP acknowledges support from the DESY computing infrastructures provided through the DESY theory group for the initial stages of this work. The work of MZ has been partly supported by the “Programma per Giovani Ricercatori Rita Levi Montalcini” granted by the Italian Ministero dell’Università e della Ricerca (MUR). We also acknowledge support from the COMETA COST Action CA22130.